A can of soda measures 12 cm tall. what is the diameter of the can if it holds 602.88 cubic cm?
Answer by checkley77(12569)

V=BASE*HEIGHT
602.88=BASE*12
BASE=602.88/12
BASE=50.24 CM^2
SEEING AS THE BASE IS A CIRCLE THEN THE AREA OF A CIRCLE=PIR^2=3.14*R^2
3.14R^2=50.24
R^2=50.24/3.14
R^2=16
R=SQRT16
R=4 CM. IS THE RADIUS OR 2*4=8 CM. IS THE DIAMETER OF THE BASE OF THE CAN.
pROOF:
12*3.14*4^2=602.88
37.68*16=602.88
602.88=602.88